OpenStudy (ashvad):

Hi guys! I really need help with this limits question.. :/ Ok, so the Question is to find the limit as x --> 16 in the function (((square root of x) -4)/(x-16)). I tried plugging in the 16, but got 0/0 (indeterminate). So, I'm trying to use the conjugate trick, but I'm getting a little lost. Can anyone help?

OpenStudy (mathmale):

First, let me verify that I have this function correct:\[(((\square \root of x) -4)/(x-16))\]

OpenStudy (mathmale):

Unfortunately, I'm unsure what you mean by "(square root of x) -4)" It would be so much better if you could present this expression symbolically. You could draw it in the Draw utility, below, or use the Equation Editor (also below). Or you could share a screen shot.

OpenStudy (ashvad):

I tried to take a screenshot of the problem! Please tell me if it works correctly, if not, I can draw it!

OpenStudy (ashvad):

\[\lim_{x \rightarrow 16} \frac{ \sqrt{x}-4 }{ x -16 }\]

OpenStudy (ashvad):

I figured out how to write it on the equation writer!

OpenStudy (alekos):

You still there?

OpenStudy (alekos):

@ashvad

OpenStudy (mathmale):

\[\lim_{x \rightarrow 16} \frac{ \sqrt{x}-4 }{ x -16 }\]

OpenStudy (mathmale):

does indeed produce 0/0 if we substitute 16 for x. This result is undefined. One approach here would be to factor the denominator as you would factor the difference between two squares. x becomes \[(\sqrt{x})^2\]

OpenStudy (mathmale):

and so the denominator is then \[(\sqrt{x}-2)(\sqrt{x}+2)\]

OpenStudy (mathmale):

\[(\sqrt{x}-4)(\sqrt{x}+4)\]

OpenStudy (mathmale):

Note my error, above; the factors involving 4 are correct, whereas those involving 2 are not.

OpenStudy (mathmale):

Now reduce the following:\[\frac{ \sqrt{x}-4 }{( \sqrt{x}-4 )(\sqrt{x}+4)}\]

OpenStudy (mathmale):

Finally, let x approach 16. Your result?

OpenStudy (ashvad):

Ohh!! I understand now! So once I do the difference of two squares factoring, I can cancel out the numerator, leaving a 1 and then when I plug in 16 for x, I end up getting 1/8 as my answer! (:

OpenStudy (ashvad):

@mathmale Thank you so much!

OpenStudy (mathmale):

1/8 sounds just right!

OpenStudy (mathmale):

My pleasure!